Guitar String

The length of the B string on a certain guitar is 59.0 cm. It vibrates at a fundamental frequency of 247.0 Hz. What is the speed of the transverse waves on the string?

If the linear mass density of the guitar string is 1.20 g/m what should the tension be when the string is in tune?

Please show all work an answer in sig-figures
Answer #1

1. Given L=0.59m, f=247.0 Hz, n=1 (fundamental)
v = 2L·f/n = 291.46 m/s

2. Given p=0.00120 kg/m
T = p·v² = 101.93871792 N

answered by: Growl